# Partial Law Invariance and Risk Measures

### v4

@@ 2.3 Partial law invariance in decision models
$\rho(X)=\max_{\mu\in\mathcal{S}}\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu=-\min_{\mu\in\mathcal{S}}\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu,$
@@
$\rho(X)=\int_{\Delta}\phi\left(\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu\right)Q(\mathrm{d}\mu)=-\int_{\Delta}\psi\left(\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu\right)Q(\mathrm{d}\mu),$
where $\phi$ is an increasing transformation, $\psi(x)=-\phi(-x)$, and $Q$ is a probability measure over $\Delta$; (c) The variational preferences model of Maccheroni et al. (2006), given by
$\rho(X)=\max_{\mu\in\Delta}\left(\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu-c(\mu)\right)=-\min_{\mu\in\Delta}\left(\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu+c(\mu)\right),$
@@ 3.1 Preferences and risk measures
Moreover, a convex risk measure is a mapping that satisfies the first three properties in the above list, and a monetary risk measure is a mapping that satisfies the first two properties in the above list; see Frittelli and Rosazza Gianin (2002) and Föllmer and Schied (2002). For background on risk measures, see Appendix A.
@@
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (3)
see Proposition 10 in Appendix A for a rigorous statement of this result. This representation is significant for two reasons. First, it links coherent risk measures to the maxmin expected utility model of Gilboa and Schmeidler (1989). With a given loss function $\ell$, the random variables $\ell(X)$ can be interpreted as a subjective loss; symmetrically, for a utility function $u$ and random wealth $X$, $u(X)$ is known as the util of $X$. Applying coherent risk measures to these subjective losses (or “dis-utils”) agrees with the maxmin expected utility model of Gilboa and Schmeidler (1989). Second, the worst-case expectation is evaluated with respect to a set of probability measures instead of finitely additive probabilities (which is the case without Fatou continuity). This distinction highlights the practical relevance of Fatou coherent risk measures, as statistical procedures yield probability measures rather than additive probabilities.
Preference relations that are represented by coherent risk measures will be called CRM preference relations, and they are described by a few properties; see Drapeau and Kupper (2013) and Proposition 12 in Appendix C. For CRM preference relations, $\mathcal{G}$-law invariance is connected to second-order stochastic dominance (SSD), which we show below. For $X,Y\in L^{\infty}$, we write $X\geq_{2}Y$ if $\mathbb{E}[\ell(X)]\geq\mathbb{E}[\ell(Y)]$ for all increasing convex functions $\ell:\mathbb{R}\to\mathbb{R}$. In the literature, the common definition is that for two random variables $Z$ and $W$ representing gains, $Z$ dominates $W$ in SSD if $\mathbb{E}[u(Z)]\geq\mathbb{E}[u(W)]$ for all increasing concave functions $u$. This is equivalent to $-W\geq_{2}-Z$, via the connection $\ell(x)=-u(-x)$, where $-W$ and $-Z$ represent losses. Next, we define SSD consistency.
@@
In the rest of the paper, we will focus on risk measures instead of preference relations. Given a monetary risk measure $\rho:L^{\infty}\to\mathbb{R}$, we call $\mathcal{A}_{\rho}=\{X\in L^{\infty}\mid\rho(X)\leq 0\}$ the acceptance set of $\rho$. Acceptance sets are a standard way to describe monetary risk measures; for some classical results regarding acceptance sets, see Appendix A. The following proposition relates partial law invariance and the acceptance set in the case of a monetary risk measure.
Proposition 5.
@@ 3.2 Representation of partially law-invariant coherent risk measures
$\mathcal{S}_{\rho}=\left\{\mu\in\mathcal{M}_{1}:\mathbb{E}^{\mu}[X]\leq\rho(X)\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ X\in L^{\infty}\right\}.$
@@
We are now ready to present a general representation for Fatou $\mathcal{G}$-law invariant coherent risk measures reminiscent of representation (3).
Theorem 1.
@@
    The risk measure $\rho$ is Fatou continuous, $\mathcal{G}$-law-invariant, and coherent;
@@
    $\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (4)
@@
We say that $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ is almost $\mathcal{G}$-law invariant if the $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closure of $\mathcal{S}$ is $\mathcal{G}$-law invariant as in parts (ii)–(iii) of Theorem 1. It is well-known that a Fatou coherent risk measure $\rho$ is law invariant if and only if its supporting set $\mathcal{S}_{\rho}\subseteq\mathcal{M}_{1}$ is law invariant; see Föllmer and Schied (2016, Theorem 4.59). Theorem 1 is reminiscent of this result, but with some subtle differences; in particular, one needs to take the restriction of elements of $\mathcal{S}$, as well as a closure of the resulting set. The first operation is clearly needed, and the second one seems essential. For a discussion of this condition and a proof of Theorem 1, see Section C.2.
Example 1.
@@
Theorem 1 also gives a feasible way to construct $\mathcal{G}$-law-invariant risk measures. We can take any law-invariant coherent risk measure $\tilde{\rho}$ on $L^{\infty}(\mathcal{G})$ (e.g., $\mathrm{ES}_{\alpha}$), and for each $\mu\in\mathcal{S}_{\tilde{\rho}}\subseteq\mathcal{M}_{1}(\mathcal{G})$ we take any set $\mathcal{S}(\mu)\subseteq\mathcal{E}(\mu)$. Setting
$\rho(X)=\sup_{\nu\in\mathcal{S}}\mathbb{E}^{\nu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ where }\mathcal{S}=\bigcup_{\mu\in\mathcal{S}_{\tilde{\rho}}}\mathcal{S}(\mu),$ (5)
@@
3.3 Two constructions
Partial law invariance of coherent risk measures naturally appear in several concrete settings, which we discuss below. The first class concerns those built on conditional expectations.
Let $\tau:L^{\infty}\to L^{\infty}(\mathcal{G})$ be a mapping such that $\tau(X)=X$ for all $X\in L^{\infty}(\mathcal{G})$; an example would be $\tau(X)=\mathbb{E}[X|\mathcal{G}]$ or, more generally, conditional convex risk measures as defined in Detlefsen and Scandolo (2005). Consider risk measures $\rho$ of the form
${\rho}(X)=\tilde{\rho}(\tau(X)),\hskip 3.0ptX\in L^{\infty},$ (6)
@@ 3.3 Two constructions
where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is a law-invariant mapping. Clearly, $\rho$ defined in (6) is $\mathcal{G}$-law invariant because $\rho(X)=\tilde{\rho}(\tau(X))=\tilde{\rho}(X)=\tilde{\rho}(Y)=\rho(Y)$ for any $X,Y\in L^{\infty}(\mathcal{G})$ with $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$.
Remark 1.

### v5

@@ 2.3 Partial law invariance in decision models
$\rho(X)=\max_{\mu\in\mathcal{S}}\int_{\Omega}\ell(X)~{}\mathrm{d}\mu=-\min_{\mu\in\mathcal{S}}\int_{\Omega}u(-X)~{}\mathrm{d}\mu,$
@@
$\rho(X)=\int_{\Delta}\phi\left(\int_{\Omega}\ell(X)~{}\mathrm{d}\mu\right)Q(\mathrm{d}\mu)=-\int_{\Delta}\widetilde{\phi}\left(\int_{\Omega}u(-X)~{}\mathrm{d}\mu\right)Q(\mathrm{d}\mu),$
where $\phi$ is an increasing transformation, $\widetilde{\phi}(x)=-\phi(-x)$, and $Q$ is a probability measure over $\Delta$; (c) The variational preferences model of Maccheroni et al. (2006), given by
$\rho(X)=\max_{\mu\in\Delta}\left(\int_{\Omega}\ell(X)~{}\mathrm{d}\mu-c(\mu)\right)=-\min_{\mu\in\Delta}\left(\int_{\Omega}u(-X)~{}\mathrm{d}\mu+c(\mu)\right),$
@@ 3.1 Preferences and risk measures
Moreover, a convex risk measure is a mapping that satisfies the first three properties in the above list, and a monetary risk measure is a mapping that satisfies the first two properties in the above list; see Frittelli and Rosazza Gianin (2002) and Föllmer and Schied (2002). For background on risk measures, see Appendix 8.
@@
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],~{}~{}~{}~{}X\in L^{\infty};$ (3)
see Proposition 8.1 in Appendix 8 for a rigorous statement of this result. This representation is significant for two reasons. First, it links coherent risk measures to the maxmin expected utility model of Gilboa and Schmeidler (1989). With a given loss function $\ell$, the random variables $\ell(X)$ can be interpreted as a subjective loss; symmetrically, for a utility function $u$ and random wealth $X$, $u(X)$ is known as the util of $X$. Applying coherent risk measures to these subjective losses (or “dis-utils”) agrees with the maxmin expected utility model of Gilboa and Schmeidler (1989). Second, the worst-case expectation is evaluated with respect to a set of probability measures instead of finitely additive probabilities (which is the case without Fatou continuity). This distinction highlights the practical relevance of Fatou coherent risk measures, as statistical procedures yield probability measures rather than additive probabilities.
Preference relations that are represented by coherent risk measures will be called CRM preference relations, and they are described by a few properties; see Drapeau and Kupper (2013) and Proposition 10.1 in Appendix 10. For CRM preference relations, $\mathcal{G}$-law invariance is connected to second-order stochastic dominance (SSD), which we show below. For $X,Y\in L^{\infty}$, we write $X\geq_{2}Y$ if $\mathbb{E}[\ell(X)]\geq\mathbb{E}[\ell(Y)]$ for all increasing convex functions $\ell:\mathbb{R}\to\mathbb{R}$. In the literature, the common definition is that for two random variables $Z$ and $W$ representing gains, $Z$ dominates $W$ in SSD if $\mathbb{E}[u(Z)]\geq\mathbb{E}[u(W)]$ for all increasing concave functions $u$. This is equivalent to $-W\geq_{2}-Z$, via the connection $\ell(x)=-u(-x)$, where $-W$ and $-Z$ represent losses. Next, we define SSD consistency.
@@
In the rest of the paper, we will focus on risk measures instead of preference relations. Given a monetary risk measure $\rho:L^{\infty}\to\mathbb{R}$, we call $\mathcal{A}_{\rho}=\{X\in L^{\infty}\mid\rho(X)\leq 0\}$ the acceptance set of $\rho$. Acceptance sets are a standard way to describe monetary risk measures; for some classical results regarding acceptance sets, see Appendix 8. The following proposition relates partial law invariance and the acceptance set in the case of a monetary risk measure.
Proposition 3.2
@@ 3.2 Representation of partially law-invariant coherent risk measures
$\mathcal{S}_{\rho}=\left\{\mu\in\mathcal{M}_{1}:\mathbb{E}^{\mu}[X]\leq\rho(X)~{}\text{for all}~{}X\in L^{\infty}\right\}.$
@@
We are now ready to present a general representation for Fatou $\mathcal{G}$-law-invariant coherent risk measures reminiscent of representation (3).
Theorem 3.3
@@
    The risk measure $\rho$ is Fatou continuous, $\mathcal{G}$-law invariant, and coherent;
@@
    $\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],~{}~{}~{}~{}X\in L^{\infty};$ (4)
@@
We say that $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ is almost $\mathcal{G}$-law invariant if the $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closure of $\mathcal{S}$ is $\mathcal{G}$-law invariant as in parts (ii)–(iii) of Theorem 3.3. It is well-known that a Fatou coherent risk measure $\rho$ is law invariant if and only if its supporting set $\mathcal{S}_{\rho}\subseteq\mathcal{M}_{1}$ is law invariant; see Föllmer and Schied (2016, Theorem 4.59). Theorem 3.3 is reminiscent of this result, but with some subtle differences; in particular, one needs to take the restriction of elements of $\mathcal{S}$, as well as a closure of the resulting set. The first operation is clearly needed, and the second one seems essential. For a discussion of this condition and a proof of Theorem 3.3, see Section 10.2.
Example 3.4
@@
Theorem 3.3 also gives a feasible way to construct $\mathcal{G}$-law-invariant risk measures. We can take any law-invariant coherent risk measure $\tilde{\rho}$ on $L^{\infty}(\mathcal{G})$ (e.g., $\mathrm{ES}_{\alpha}$), and for each $\mu\in\mathcal{S}_{\tilde{\rho}}\subseteq\mathcal{M}_{1}(\mathcal{G})$ we take any set $\mathcal{S}(\mu)\subseteq\mathcal{E}(\mu)$. Setting
$\rho(X)=\sup_{\nu\in\mathcal{S}}\mathbb{E}^{\nu}[X],~{}~{}X\in L^{\infty},\mbox{~{}~{}where }\mathcal{S}=\bigcup_{\mu\in\mathcal{S}_{\tilde{\rho}}}\mathcal{S}(\mu),$ (5)
@@
4 Multiple sources
A natural corollary of Theorem 3.3 characterizes all Fatou $\mathfrak{G}$-law invariant coherent risk measures.
Theorem 4.1
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],~{}~{}~{}~{}X\in L^{\infty}$
Appendix 11 contains an example of a $\mathfrak{G}$-law-invariant coherent risk measure given by $\rho(X)=\sup_{\mathcal{G}\in\mathfrak{G}}\rho_{\mathcal{G}}(\mathbb{E}[X|\mathcal{G}])$, where each $\rho_{\mathcal{G}}$ is law invariant and coherent, along with its representation in the form of Theorem 4.1.
  1. (a)
  2. (b)
It is clear that both (a) and (b) are stronger than $\mathfrak{G}$-law invariance, and moreover, (b) is stronger than (a). In Example 4.2 below, we illustrate a situation in which (b) is not desirable, whereas (a) may be reasonable.
Example 4.2
$\rho(X)=\sup_{\lambda\in[0,1]}\mathrm{ES}^{\mu_{\lambda}}_{\alpha}(X),~{}~{}~{}X\in L^{\infty}.$
Our current formulation of $\mathfrak{G}$-law invariance has an advantage over property (a), that is, the flexibility for modeling the situation where different sources are associated with different levels of model uncertainty (ambiguity), illustrated by Example 4.3 below.
Example 4.3
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}\left(\mathbb{E}^{\mu}[\ell(X)]-\frac{1}{\theta_{i}}R(\mu,\mathbb{P})\right)=\frac{1}{\theta_{i}}\log\mathbb{E}\left[e^{\theta_{i}\ell(X)}\right],~{}~{}X\in L^{\infty}(\mathcal{G}_{i}),$ (6)
In Example 4.3, even if model uncertainty exists in each source $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$, the risk measure is law invariant on each subspace as a result of robustification, discussed by Hansen and Sargent (2001).
Example 4.4
[…]
5 Extensions
In this section, we present three extensions of partially law-invariant coherent risk measures. First, in Section 5.1, we introduce a technical tool called coherent adjustments, which admit interesting mathematical properties. Coherent adjustments allow us to reformulate Theorem 3.3 and serve as the foundation for the subsequent two extensions. Second, in Section 5.2, we strengthen partial law invariance to a property, which still generalizes law invariance, that allows us to extend the representation of Kusuoka (2001). Finally, in Section 5.3, we extend some of the results regarding partially law-invariant coherent risk measures to partially law-invariant convex risk measures.
5.1 Coherent adjustments and an alternative representation
Definition 5.1
Let $\mu\in\mathcal{M}_{1}(\mathcal{G})$. From the above properties, it is straightforward that the supremum of $\mu$-coherent adjustments is again a $\mu$-coherent adjustment. A simple example of a Fatou $\mu$-coherent adjustment is $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ for any $\nu\in\mathcal{E}(\mu)$, which is easy to check. In Section 12.1, Proposition 12.2, shows that, indeed, the supremum of the form $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ contains all possible Fatou $\mu$-coherent adjustments.
Remark 5.2
The name “$\mu$-coherent adjustment” is explained by Proposition 5.3.
Proposition 5.3
  1. (i)
  2. (ii)
If we assume additivity on $L^{\infty}(\mathcal{G})$ and $\mathcal{G}$-law invariance for a Fatou coherent risk measure $\rho$, then $\rho$ has a form $\mathbb{E}^{\mathbb{P}}+\tau$ for $\tau\in\mathrm{CA}(\mathbb{P}|_{\mathcal{G}})$ that is similar to Proposition 5.3 (ii); see Section 12.2.
For some $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, by Proposition 5.3, the following risk measure
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left(\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right),~{}~{}~{}X\in L^{\infty},$ (7)
is coherent, as it is the supremum of coherent risk measures. Theorem 5.4 below shows that this mapping is Fatou continuous and that all Fatou coherent risk measures have this representation.
To explain why $\mu$-coherent adjustments help us extend a coherent risk measure on $L^{\infty}(\mathcal{G})$ to $L^{\infty}$, let $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ be a law-invariant coherent risk measure. For a convex $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-dense in $\mathcal{S}_{\tilde{\rho}}$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, let $\rho$ be given by (7). It is easy to see that $\rho|_{L^{\infty}(\mathcal{G})}=\tilde{\rho}$, meaning that it is $\mathcal{G}$-law invariant.
Theorem 5.4
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left(\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right),~{}~{}~{}X\in L^{\infty}.$ (8)
Remark 5.5
There is a straightforward way to switch between representations (4) and (8) for a Fatou and $\mathcal{G}$-law-invariant coherent risk measure. Given a convex set $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that represents $\rho$ in Theorem 3.3, define $\mathcal{R}=\mathcal{S}^{\mathcal{G}}$, which is almost $\mathcal{G}$-law invariant. For each $\mu\in\mathcal{R}$, define $\mathcal{S}(\mu)=\mathcal{E}(\mu)\cap\mathcal{S}$ and $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}(\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X])$. These can be used in representation (8). Conversely, let an almost $\mathcal{G}$-law invariant convex set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ representing $\rho$ in Theorem 5.4 be given. For $\mu\in\mathcal{R}$, let $\mathcal{S}_{\tau_{\mu}}$ be the $w(\mathcal{M},L^{\infty})$-closed and convex set in Proposition 12.2 (see Section 12.1) that represents $\tau_{\mu}$. It is easy to show that $\mathcal{S}=\bigcup_{\mu\in\mathcal{R}}\mathcal{S}_{\tau_{\mu}}$ is convex and $S^{\mathcal{G}}=\mathcal{R}$. We can use $\mathcal{S}$ in representation (4). The above connection clarifies the construction in (5).
Uniqueness of representation (8) under a stronger notion of continuity is addressed in Section 12.3. Representation (8) has little resemblance to the classical representation result of Kusuoka (2001) for law-invariant coherent risk measures. The following section addresses this gap by further imposing a necessary and sufficient condition for a similar representation along that direction.
5.2 Strong partial law invariance and Kusuoka-type representation
Next, we introduce a new property, which is stronger than partial law invariance and is necessary and sufficient to reduce (8) to a Kusuoka-type representation; see Proposition 8.3 in Appendix 8.
$\rho(Z+X)=\rho(Z+Y).$
Lemma 5.6
@@ 5.2 Strong partial law invariance and Kusuoka-type representation
The mapping $\rho$ is strongly $\mathcal{G}$-law invariant if and only if for all $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$ with $Y\overset{\mathrm{d}}{=}_{\mathbb{P}}\mathbb{E}[X|\mathcal{G}]$, we have
$\rho(X)=\rho(X-\mathbb{E}[X|\mathcal{G}]+Y).$
An example of strongly $\mathcal{G}$-law invariant risk measure $\rho$ is $\rho(X)=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$, where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is law invariant. Examples of functionals satisfying $\mathcal{G}$-law invariance but not strong $\mathcal{G}$-law invariance will be provided in Section 12.5.
[…]
$\sup_{\mu\in\mathcal{M}_{Q}}\mathbb{E}^{\mu}[X]=\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])~{}Q(\mathrm{d}\alpha);$ (10)
Theorem 5.7
Let $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
  2. (ii)
    $\rho(X)=\sup_{Q\in\mathcal{D}}\left(\int_{[0,1)}\mathrm{ES}_{\alpha}\left(\mathbb{E}[X|\mathcal{G}]\right)~{}Q(\mathrm{d}\alpha)+\tau_{Q}(X)\right).$ (11)
$\rho(X)=\sup_{Q\in\mathcal{D}}\int_{[0,1)}\mathrm{ES}_{\alpha}(X)~{}Q(\mathrm{d}\alpha),\hskip 3.0ptX\in L^{\infty},$
because in this case $\mathbb{E}[X|\mathcal{G}]=X$, $\tau_{Q}(X)=0$, and strong $\mathcal{G}$-law invariance is law invariance. A simple example is to choose $\tau_{Q}=0$ for all $Q\in\mathcal{D}$, which yields the risk measure $\rho:X\mapsto\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$ studied in Proposition 6.5 below.
5.3 Convex risk measures
This section discusses convex risk measures and their relation to the variational preferences model of Maccheroni et al. (2006). We provide representations for partially law-invariant convex risk measures akin to Theorem 5.4. A risk measure $\rho:L^{\infty}\to\mathbb{R}$ is Fatou and convex if and only if there exists $\alpha:\mathcal{M}_{1}\to(-\infty,\infty]$ such that
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}(\mathbb{E}^{\mu}[X]-\alpha(\mu)),~{}~{}~{}X\in L^{\infty};$
Similar to CA-assignments, we say that $(\tau_{\mu})_{\mu\in\mathcal{M}_{1}(\mathcal{G})}$ is a CoA-assignment if for all $\mu\in\mathcal{M}_{1}(\mathcal{G})$, $\tau_{\mu}$ is a Fatou $\mu$-convex adjustment and the mapping $\mu\mapsto\tau_{\mu}(0)$ is concave. The following result is the convex version of Theorem 5.4.
Theorem 5.8
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left(\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right)~{}~{}~{}X\in L^{\infty}.$
The risk measures ER in Example 4.4 are the most common examples of convex risk measures. We next discuss a few $\mathcal{G}$-law-invariant versions of ER, which shed some light on different ways to construct partially law-invariant convex risk measures. In the examples below, partial law invariance is not necessarily associated with the interpretation of model uncertainty.
$\mathrm{ER}_{\beta}(X|\mathcal{G})=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right],~{}~{}~{}X\in L^{\infty}.$
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathbb{E}[\mathrm{ER}_{\beta}(X|\mathcal{G})]=\mathbb{E}\left[\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right],~{}~{}~{}X\in L^{\infty}.$
The second is to compute the ER of the conditional mean, that is,
$\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathrm{ER}_{\beta}(\mathbb{E}[X|\mathcal{G}])=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta\mathbb{E}[X|\mathcal{G}]}\right],~{}~{}~{}X\in L^{\infty}.$
$\mathrm{ER}_{\beta}(\mathrm{ER}_{\beta}(X|\mathcal{G}))=\frac{1}{\beta}\log\mathbb{E}\left[\exp\left(\beta\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right)\right]=\mathrm{ER}_{\beta}(X),~{}~{}~{}X\in L^{\infty}.$
[…]
if $Y$ is independent of $\mathcal{G}$, then
[…]
6 Explicit classes and applications
In this section, we propose a few explicit constructions of partially law-invariant risk measures and illustrate with numerical examples. Recall our notation ()=$\mathcal{E}(\mathbb{P})=\{\nu\in\mathcal{M}_{1}:\nu|_{\mathcal{G}}=\mathbb{P}\}$, which will be used frequently below.
6.1 Partially law-invariant ES
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X),~{}~{}~{}~{}X\in L^{\infty},$
where $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$ and satisfies some regularity (to be made rigorous soon). Afterward, we numerically implement the results of this section and see how decisions in a simple investment vary as we change the parameters in the risk measure.
We focus on ES here for two reasons. First, ES is the standard regulatory risk measure in the banking sector (BCBS (2019)), and hence the most relevant for financial applications. Second, ES has special advantages in optimization (Rockafellar and Uryasev (2002)), and the partially law-invariant version of ES admits tractable formulas (Proposition 6.3 below).
Definition 6.1
  1. (i)
    For all $X\in\mathcal{X}$, $X\leq Y$;
  2. (ii)
    if $Z\in L^{\infty}$ and $X\leq Z$ for all $X\in\mathcal{X}$, then $Y\leq Z$.
Definition 6.2
Given certain conditions on $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$, the $\mathcal{G}$-law invariant risk measure $\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}$ mentioned in Section 2.2 can be determined via a minimization formula, similar to the formulation of ES by Rockafellar and Uryasev (2002). Below, a coherent conditional risk measure is a mapping from $L^{\infty}$ to $L^{\infty}(\mathcal{G})$ satisfying the four axioms of coherence but with constants replaced by elements of $L^{\infty}(\mathcal{G})$; see Detlefsen and Scandolo (2005).
Proposition 6.3
$\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}\left[\rho_{\mathcal{G}}^{\mathcal{Q}}\left((X-x)^{+}\right)\right]\right),~{}~{}~{}X\in L^{\infty},$
$\rho_{\mathcal{G}}^{\mathcal{Q}}(X)=\operatorname*{ess\text{-}sup}_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X|\mathcal{G}],~{}~{}~{}~{}X\in L^{\infty}.$
Moreover, the set $\mathcal{Q}_{\beta}=\left\{\mu\in\mathcal{E}(\mathbb{P}):~{}D_{\mu}\leq 1/(1-\beta)\right\}$ for $\beta\in[0,1)$ is $w(\mathcal{M},L^{\infty})$-compact and $\mathcal{G}$-convex.
Denote by $\mathrm{ES}_{\beta}(X|\mathcal{G})=\rho_{\mathcal{G}}^{\mathcal{Q}_{\beta}}(X)$, $X\in L^{\infty}$, which is the conditional version of ES in Detlefsen and Scandolo (2005). The benefit of using $\mathcal{Q}_{\beta}$ is that under tractable assumptions of the conditional distribution of $X$ given $\mathcal{G}$, $\mathbb{E}[\mathrm{ES}_{\beta}((X-x)^{+}|\mathcal{G})]$ can be computed using conditional distributions. Section 6.3 offers a numerical illustration of this idea.
6.2 Two constructions of partially law-invariant coherent risk measures
We next present two approaches to constructing partially law-invariant coherent risk measures. The first class arises from mappings analogous to conditional expectations.
Let $\psi:L^{\infty}\to L^{\infty}(\mathcal{G})$ be a mapping such that $\psi(X)=X$ for all $X\in L^{\infty}(\mathcal{G})$; an example would be $\psi(X)=\mathbb{E}[X|\mathcal{G}]$ or, more generally, the coherent conditional risk measures. Consider risk measures $\rho$ of the form
${\rho}(X)=\tilde{\rho}(\psi(X)),\hskip 3.0ptX\in L^{\infty},$ (12)
where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is a law-invariant mapping. Clearly, $\rho$ defined in (12) is $\mathcal{G}$-law invariant because $\rho(X)=\tilde{\rho}(\psi(X))=\tilde{\rho}(X)=\tilde{\rho}(Y)=\rho(Y)$ for any $X,Y\in L^{\infty}(\mathcal{G})$ with $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$.
Remark 6.4